. The red spectral line of lithium occurs at 671nm.calculate the energy of one photon.


1
Expert's answer
2022-02-28T11:02:42-0500

The frequency of light can be found using c = νλ by rearranging to give ν = c/λ = (3.00 x 108 m/s) ÷ (6.71 x 10-7 m) = 4.47 x 1014/s.

The frequency of this light is

 ν = C/ λ = 3.00 x 108 m/s / 6.71x 10-7 m = 4.47 x 1014/s

Hence, the energy of one photon is

  E = hν = 6.63 x 10-34 J.s x 4.47 x 1014/s = 2.96 x 10-19 J